# Smriti Singh*1, Shuvam Keshari*1, Vinija Jain2, Aman Chadha2,3

### v3

@@ Abstract
Socioeconomic bias in society exacerbates disparities, influencing access to opportunities and resources based on individuals’ economic and social backgrounds. This pervasive issue perpetuates systemic inequalities, hindering the pursuit of inclusive progress as a society. In this paper, we investigate the presence of socioeconomic bias, if any, in large language models. To this end, we introduce a novel dataset SilverSpoon, consisting of 3000 samples that illustrate hypothetical scenarios that involve underprivileged people performing ethically ambiguous actions due to their circumstances, and ask whether the action is ethically justified. Further, this dataset has a dual-labeling scheme and has been annotated by people belonging to both ends of the socioeconomic spectrum. Using SilverSpoon, we evaluate the degree of socioeconomic bias expressed in large language models and the variation of this degree as a function of model size. We also perform qualitative analysis to analyze the nature of this bias. Our analysis reveals that while humans disagree on which situations require empathy toward the underprivileged, most large language models are unable to empathize with the socioeconomically underprivileged regardless of the situation. To foster further research in this domain, we make SilverSpoon and our evaluation harness publicly available [^1] https://github.com/keshariS/socioEcoBias.
@@ 1 Introduction
Bias in large language models (LLMs) continues to be a pervasive issue, and a considerable amount of research has been done in this area. While gender bias Kotek et al. (2023); Garrido-Muñoz et al. (2023); Gallegos et al. (2023), racial bias Omiye et al. (2023); Zack et al. (2024) religion-based bias Abid et al. (2021), and other kinds of bias Liang et al. (2021); Nadeem et al. (2020); Kirk et al. (2021); Khandelwal et al. (2023) in these models have been investigated, one of the under-explored key dimensions along which these models may be biased is socioeconomic bias. To the best of our knowledge, there has been no research analyzing whether LLMs exhibit socioeconomic bias. In this work, we ask the question, are large language models perpetrators of socioeconomic bias? Our question is also motivated by the fact that these models are typically trained on very large amounts of data taken from the internet, and internet access itself is usually a reflection of at least some socioeconomic privilege Powell et al. (2010). Further, opinions on the internet themselves may be reflective of biases against certain socioeconomic groups, which these models might have unintentionally picked up.
@@
To ground our analysis, we present SilverSpoon, a dataset consisting of 3000 questions about socioeconomically underprivileged people facing challenging dilemmas. These questions have been generated by a combination of thoughtful prompting of GPT4 and text augmentation techniques. answered by annotators belonging to both ends of the socioeconomic spectrum, and these answers are considered gold labels for this study. This is, to the best of our knowledge, the first dataset that contains high-quality labels in the form of answers for questions meant to analyze how perception changes across the socioeconomic spectrum. Through data annotation, prompt engineering, and qualitative analysis, we aim to answer the following research questions:
@@
  1. RQ1: Do large language models exhibit socioeconomic bias when tasked with understanding the challenging conditions of socioeconomic struggle?
  2. RQ2: How often do these models express the opinion of the socioeconomically privileged?
  3. RQ3: Are LLMs able to exhibit empathy for the socioeconomically underprivileged in difficult situations?
  4. RQ4: Do LLMs of varying sizes express varying degrees of socioeconomic bias or empathy toward the underprivileged?
To this end, our research work makes the following contributions:
Our Contributions We present the first dual-label dataset consisting of 3k samples, designed to analyze ethical differences on the opposing ends of the socioeconomic spectrum. Notably, we provide labels from annotators that belong to both ends of this spectrum. We present, to the best of our knowledge, the first analysis of socioeconomic bias in state-of-the-art large language models. We provide an extensive quantitative and qualitative analysis of the variation of responses given by differing SoTA LLMs.
We find that almost all LLMs that we test are unable to exhibit any empathy toward the socioeconomically underprivileged. This is concerning, because it may have unintended ill effects on several downstream applications, especially those which deal with fields already impacted by socioeconomic bias. Notably, Alpaca exhibits high disagreement with the gold labels obtained from the lower end of the socio-economic spectrum, and GPT-4 is surprisingly unempathetic toward the underprivileged (even compared to other LLMs). An overview of our entire pipeline is demonstrated in Figure 2.
@@ 2. RQ2: How often do these models express the opinion of the socioeconomically privileged?
Figure 2: A step-by-step illustration of our pipeline demonstrating the three major components as we look to answer the question “Are the underprivileged legally wrong… or morally right?"
@@ 2 The SilverSpoon Dataset
Here, we present SilverSpoon, a dataset that consists of questions that examine hypothetical harsh actions underprivileged people take under extreme circumstances, along with gold-label annotations to indicate whether these actions are wrong. Notably, these annotations are provided by people from both ends of the socioeconomic spectrum. Figure 1 provides a visualization of our dataset. Figure 4 illustrates some examples of the structure of this dataset. (This dataset is under license CC BY 4.0)
@@ 2.1 Dataset Design
The design of this dataset is ethically motivated by the question, "Does socioeconomic status (SES) contribute to the ability to empathize with socioeconomically underprivileged people?". Many studies from psychology and sociology show that SES may be a factor in the ability to experience empathy Liu et al. (2023); Pishghadam et al. (2022); Jolliffe and Farrington (2006); Li and Siu (2021); Love (2020). Further, the consensus from social media seems to be that the people on the higher end of the socioeconomic spectrum don’t understand the challenges that the people on the opposite end of the spectrum face  [^2] Reddit Thread 1, Reddit Thread 2, Reddit Thread 3. However, the answer to this question, and the effect it may have remains unanswered as far as LLMs are concerned.
Thus, for the design of this dataset, we carefully craft a prompt that consists of hypothetical harsh actions that socioeconomically underprivileged people may take, and ask GPT4 to generate situations that are similar to this. Examples of the result of this prompt can be observed above, in Figure 4. We ask GPT4 to generate 1000 such samples. The prompt we use for this task is available in Appendix E. Then, we use text augmentation to perform synonym replacement on each of these 1000 samples, with an empirically obtained similarity threshold of $0.8$. Thus, our dataset consists of 3000 samples. By adopting this approach in our dataset design, we save computational resources in data generation and human labor in the annotation process.
@@
2.2 Annotation
To annotate this data, we had four annotators answer the questions in the first 1000 samples. Of the four annotators, two annotators were brought up in families that can be categorized as belonging to the upper end of the socioeconomic spectrum (referred to as the upper-end category from here on), while two annotators were brought up in families that can be categorized as belonging to the lower end of the socioeconomic spectrum (referred to the lower end category from here on). The upper-end label annotators have an income greater than 200K USD per annum. The lower-end label annotators have an income of less than 40K USD per annum.
The annotation task was described as answering a set of ethical questions in which either P1 (the underprivileged person) or the "other person" is wrong. The exact instruction set provided to annotators can be found in the Appendix B. To create the gold labels, we took the intersection of their annotations and randomly sampled where there was a disagreement between two annotators that belong to the same category. We found that across categories, the annotators had a Cohen Kappa score of 0.141. Within categories, the Cohen Kappa scores were 0.781 and 0.793 for the upper-end and lower-end labels respectively. Interestingly, we found that 46.67% of the labels from the upper-end label support the idea that the "other person" is wrong, while this number is 36.30% for the lower-end label.
@@ 2.2 Annotation
2.3 Data Analysis
Figure 5: SilverSpoon data distribution. Please refer to Section 2.3 for information about each topic.
The average length of the samples in our dataset is about 15 words per sample. To provide some insight into the most common issues investigated in this dataset, we employ LDA, a common topic modeling technique. The top 10 topics in our dataset are illustrated in Figure 1. We describe these topics by listing the top 5 most frequently occurring words respectively in the Appendix A. Concurrent with trends in research Phelan et al. (1995); Kraus et al. (2012); Manstead (2018), our dataset has an emphasis on the issues of theft, accessing school, and homelessness. Figure 3 illustrates samples from each of these issues. Further, Figure 5 displays the distribution of the top 10 topics in our dataset. We also illustrate the distribution of the upper-end labels and lower-end labels in these topics. This is shown in Figures 6 and 7.
@@ 3 Study Design
We experimented with GPT-4 Bubeck et al. (2023), GPT-3.5 Brown et al. (2020), Llama2-Chat-13B Touvron et al. (2023), Llama2-Chat-7B, Alpaca-7B Taori et al. (2023), Zephyr-7B Tunstall et al. (2023), and Yi-6B 01.ai (2024). Our main prompt asks the model to decide who is wrong in the given scenario, either "P1" or the "other person", for each of the 3000 samples in our dataset. Further, to perform high-caliber qualitative analysis, we elicit concise natural language explanations that justify the models’ answers. Notably, we find that the prompt used for GPT4 does not allow other models to engage with the text properly. Thus, we have a different prompt for the other models. We illustrate both prompts in Appendix E and describe the results of our prompting in Section 4.
We choose to perform the prompting in a zero-shot manner because we are interested in examining the responses of models without any examples being provided to them. We hypothesize that even 1 example could introduce extraneous bias. Our experiments are run using A100s available on Google Colab and take a total of approximately 69 hours.
@@ 4 Results
4.1 Quantitative Analysis
We report per class accuracy, Cohen’s Kappa, Jaccard Similarity, Dice Similarity, and Matthew’s Correlation Coefficients against the upper-end labels and lower-end labels respectively. We do not use F1, precision, and recall in our evaluation since these metrics are calculated against a true-positive class, and our goal is to understand the similarity between model responses and each category of labels. We present multiple metrics in order to illustrate a complete picture of our quantitative analysis. Table 1 presents these results, To convert the model responses to a binary scale, we employ automatic evaluation, asking GPT4 to assess various responses and assign them to Class 1: Supporting the socioeconomically underprivileged, or Class 0: Otherwise. The prompt for this can be found in the appendix E.6.
@@ 4.1 Quantitative Analysis
The range of values for accuracy, JSC, and DSC is 0 to 1 whereas the range for CKS and MCC is -1 to 1. For the latter two metrics, 0 indicates random chance agreement. We find that Alpaca has a negative CKC and MCC implying that its correlation with the lower-end labels is less than random, hence it agrees more with the upper class labels. AppendixD offers a detailed discourse on the metrics of our choice and what they signify.
@@
4.2 Qualitative Analysis
4.2.1 GPT
We observe that GPT-4 and GPT-3.5 tend to answer "P1 is wrong". In the rare occurrence that either of these models says other, the explanation they offer almost always has to do with the fact that P1 is in immediately life-threatening environments. For example, these models state that P1 is wrong when they cannot access education, or even when they cannot access medicine, but can sympathize with P1 when they break into shelters to survive "brutally cold winters". Figure 9 illustrates how GPT4 seems to consider only the legal aspect of these situations, without any perceivable concept of human concern or empathy. As seen in Figure 17, we find that GPT-4 exhibits a complete lack of empathy when it comes to the socioeconomically underprivileged, even when sensitive issues like domestic abuse are involved.
As far as GPT-3.5 is concerned, we observe that it does tend to empathize with the underprivileged more than GPT-4, by 19.33% which can be seen in Figure 8. Some examples of this are presented in Figure 10. The topic-wise distribution of GPT3.5s responses can be found in Figure 18.
Figure 10: Detox is wrong but stealing is OK? Examples of unusual responses by GPT-3.5.
4.2.2 Llama-2-Chat
Figure 11: Both are wrong? Examples of unusual responses by Llama-13B.
We find that Llama-2-Chat-13B says "both are wrong" more often than other models. (For quantitative evaluation, we discard these responses.) Llama-2-Chat-13B has a very similar response distribution to GPT-3.5, a fact that is reflected in Figure 8. The most frequently occurring explanation this model offers for siding with the underprivileged is along the lines of "Respecting others’ boundaries" is important. In this manner, Llama-2-Chat displays more sympathy than GPT-4 (but marginally lesser than GPT-3.5) toward the socioeconomically underprivileged. However, it is interesting to see how Llama2 reacts to different situations, and a few examples have been demonstrated in Figure 11. Figure 19 shows a topic-wise analysis of model responses.
Llama-2-Chat-7B, on the other hand, despite having a response distribution that is closest to GPT-4 (as seen in Figure 20), stands out for it’s illogical responses. Some of them are highlighted in Figure 12. One possible cause for this behavior concerning this problem set is that the model size is much smaller than the others we test.
Figure 12: You are wrong for even asking this? Examples of unusual responses by Llama-7B. (L.E. stands for lower-end labels and U.E. stands for upper-end labels.)
4.2.3 Alpaca
In our observation, Alpaca frequently fails at explaining its responses, despite being explicitly asked for one. When it does offer explanations, it is almost always for the response "P1 is wrong". Like Llama-2-Chat-7B, it also answers "both" and "neither" to the question, "Who is wrong?". Refer to Figure 13 for examples and Figure 21 for label distribution analysis.
4.2.4 Zephyr
We find that Zephyr is the most coherent of the 7B models that we test on. Unlike Alpaca-7B and Llama-2-Chat-7B, this model almost always provides accurate explanations for its answer. Even though it is not always able to exactly follow the instructions of the prompt and answer only "p1 is wrong" or "other is wrong", it can justify its answer understandably. Examples are illustrated in Figure 14 and response distribution per topic is shown in Figure 22.
4.2.5 Yi

### v4

# Smriti Singh*1, Shuvam Keshari*1, Vinija Jain2  , Aman Chadha3,4

@@ Abstract
Socioeconomic bias in society exacerbates disparities, influencing access to opportunities and resources based on individuals’ economic and social backgrounds. This pervasive issue perpetuates systemic inequalities, hindering the pursuit of inclusive progress as a society. In this paper, we investigate the presence of socioeconomic bias in large language models. To this end, we introduce a novel dataset SilverSpoon, consisting of 12000 samples that provide a multifaceted analysis of this complex issue. This dataset has three subsets. The first 3000 samples focus on normative judgement evaluation, consisting of hypothetical scenarios in which people of different socioeonomic class make difficult decisions. This subset of the dataset has a dual-labeling scheme and has been annotated by people belonging to both ends of the socioeconomic spectrum. The second subset of this dataset focuses on demographic driven profession prediction, and consists of 8000 samples that investigate socioeconomic bias across a plethora of combinations of gender, race and location. Finally, the third subset of the dataset focuses on contextual narrative bias analysis. This subset consists of 1000 LLM generated stories, which have been leveraged to detect the presence of subtle stereotypes against certain socioeconomic classes belonging to various demographic groups. Using SilverSpoon, we evaluate the degree of socioeconomic bias expressed in state-of-the-art large language models. We also perform extensive quantitative and qualitative analysis to analyze the nature of this bias. Our analysis reveals that state-of-the-art large language models exhibit implicit and explicit socioeconomic bias, which is further augmented by stereotypes emanating from a combination of gender bias and racial bias. To foster further research in this domain, we make SilverSpoon and our evaluation harness publicly available.
@@ 1 Introduction
Bias in large language models (LLMs) continues to be a pervasive issue, and a considerable amount of research has been done in this area. While gender bias Kotek et al. (2023); Garrido-Muñoz et al. (2023); Gallegos et al. (2023), racial bias Omiye et al. (2023); Zack et al. (2024) religion-based bias Abid et al. (2021), and other kinds of bias Liang et al. (2021); Nadeem et al. (2020); Kirk et al. (2021); Khandelwal et al. (2023) in these models have been investigated, one of the under-explored key dimensions along which these models may be biased is socioeconomic bias. To the best of our knowledge, there has been just one research paper analyzing whether LLMs exhibit socioeconomic biasArzaghi et al. (2024). While their paper offers valuable insight about intrinsic socioeconomic bias, we approach this issue in a more comprehensive and holistic manner, including demographic driven bias analysis, but expanding to more nuanced issues, like normative judgement and contextual narrative bias. In this work, we ask the question, are large language models perpetrators of socioeconomic bias? Our question is also motivated by the fact that these models are typically trained on very large amounts of data taken from the internet, and internet access itself is usually a reflection of at least some socioeconomic privilege Powell et al. (2010). Further, opinions on the internet themselves may be reflective of biases against certain socioeconomic groups, which these models might have unintentionally picked up.
@@
To ground our analysis, we present SilverSpoon, a dataset consisting of 12000 samples. Of these, there are 3000 questions about socioeconomically underprivileged people facing challenging dilemmas. These questions have been generated by a combination of thoughtful prompting of GPT4 and text augmentation techniques. answered by annotators belonging to both ends of the socioeconomic spectrum, and these answers are considered gold labels for this study. This is, to the best of our knowledge, the first dataset that contains high-quality labels in the form of answers for questions meant to analyze how perception changes across the socioeconomic spectrum. The dataset further consists of 8000 combinations of names and location across race and gender, which we leverage to prompt SOTA LLMs and uncover potential biases linked to different demographic groups through the task of profession prediction, to analyze how socioeconomic bias in LLMs varies across demographic groups. SilverSpoon also consists of 1000 story generation prompts which allow for a deeper analysis of contextual narrative socioeconomic biases in the portrayal of lifestyles, challenges and social status. Through data annotation, prompt engineering, and qualitative analysis, we aim to answer the following research questions:
@@
  1. RQ1: To what extent do large language models exhibit socioeconomic bias when tasked with understanding or empathizing with the experiences of the socioeconomically underprivileged, particularly in challenging situations?
  2. RQ2: How do large language models of varying sizes express opinions or biases related to socioeconomic privilege, and does model size impact their ability to demonstrate empathy toward the underprivileged?
  3. RQ3: To what extent do large language models exhibit biases in profession prediction based on name and location, and how are these biases influenced by race, gender, and socioeconomic status?
  4. RQ4: How do large language models implicitly portray socioeconomic status and privilege in generated narratives based on demographic cues such as name and location, and what underlying biases emerge from these portrayals?
We find that most LLMs are unable to exhibit any empathy for socioeconomically underprivileged people in difficult conditions, and that existing socioeconomic bias is further augmented by stereotypical beliefs.
@@
Figure 2: A step-by-step illustration of our pipeline demonstrating the three major components as we look to answer our research questions.
@@ 2 The SilverSpoon Dataset
Here, we present SilverSpoon, a dataset that presents a multifacted analysis of socioeconomic bias in SoTA LLMs. Given the complex nature of socioeconomic bias, we consider the following while designing this dataset: First, we are interested in evaluating normative judgements. Asking questions that instruct SoTA LLMs to judge actions or choices based on socieocoonmic status allows us to analyze how these models endorse or challenge societal norms and values related to privilege, wealth and social inequality. Second, we are interested in demographic driven profession prediction as a method for analyzing socioeconomic bias because it provides an opportunity to investigate implicit stereotypical associations these language models may make between various demographic cues and occupational outcomes. Our hypothesis is that this approach will also help highlight patterns of bias linked to these demographic groups, if they exist. Finally, we are interested in contextual narrative bias analysis, which we hypothesize will uncover more subtle stereotypes and assumptions of social class and privilege, by examining how these language models construct narratives around individuals based on demographic cues. Figure 1 provides a visualization of our dataset. Figure 4 illustrates some examples of the structure of this dataset. (This dataset is under license CC BY 4.0)
@@ 2.1 Dataset Design
2.1.1 Part 1: Normative Judgement Evaluation
The design of this subset of our dataset is ethically motivated by the question, "Does socioeconomic status (SES) contribute to the ability to empathize with socioeconomically underprivileged people?". Many studies from psychology and sociology show that SES may be a factor in the ability to experience empathy Liu et al. (2023); Pishghadam et al. (2022); Jolliffe and Farrington (2006); Li and Siu (2021); Love (2020). Further, the consensus from social media seems to be that the people on the higher end of the socioeconomic spectrum don’t understand the challenges that the people on the opposite end of the spectrum face [^1] Reddit Thread 1, Reddit Thread 2, Reddit Thread 3. However, the answer to this question, and the effect it may have remains unanswered as far as LLMs are concerned.
Thus, for the design of this part of the dataset, we carefully craft a prompt that consists of hypothetical harsh actions that socioeconomically underprivileged people may take, and ask GPT4 to generate situations that are similar to this. Examples of the result of this prompt can be observed above, in Figure 4. We ask GPT4 to generate 1000 such samples. The prompt we use for this task is available in Appendix K. Then, we use text augmentation to perform synonym replacement on each of these 1000 samples, with an empirically obtained similarity threshold of $0.8$. Thus, this subset of our dataset consists of 3000 samples. By adopting this approach in our dataset design, we save computational resources in data generation and human labor in the annotation process.
@@ 2.1.1 Part 1: Normative Judgement Evaluation
To annotate this data, we had six annotators answer the questions in the first 1000 samples. Of the six annotators, three annotators were brought up in families that can be categorized as belonging to the upper end of the socioeconomic spectrum (referred to as the upper-end category from here on), while three annotators were brought up in families that can be categorized as belonging to the lower end of the socioeconomic spectrum (referred to the lower end category from here on). The upper-end label annotators have an income greater than 200K USD per annum. The lower-end label annotators have an income of less than 40K USD per annum.
The annotation task was described as answering a set of ethical questions in which either P1 (the underprivileged person) or the "other person" is wrong. The exact instruction set provided to annotators can be found in the AppendixB. To create the gold labels, we took the intersection of their annotations and considered a majority vote where there was a disagreement between two annotators that belong to the same category. We found that across categories, the annotators had a Cohen Kappa score of 0.141. Within categories, the Cohen Kappa scores were 0.781 and 0.793 for the upper-end and lower-end labels respectively. Interestingly, we found that 46.67% of the labels from the upper-end label support the idea that the "other person" is wrong, while this number is 36.30% for the lower-end label. A comprehensive dataset analysis can be found in AppendixC.
@@
2.1.2 Part 2: Demographic Driven Profession Prediction
In this study, we constructed a dataset by selecting the 50 most common names for both men and women from four distinct ethnic groups: Black, White, Hispanic, and Indian. These names were then combined with the 10 richest and 10 poorest cities in the United States to create a comprehensive set of demographic-based prompts. This approach generated a wide range of combinations, representing diverse socioeconomic and racial backgrounds. For now, we limit our analysis to locations within one country to minimize the error of not recognizing the effect of various cultural practices/beliefs around the world. A full list of the resources we used to compile this information together can be found in AppendixF.
For example, if we take into consideration the names John, Meera, Shaw and Gloriana paired with the cities Los Angeles and Detroit, our generation method results in a dataset that looks like this:
Name Gender Race Location
John Male White Los Angeles
John Male White Detroit
Meera Female Indian Los Angeles
Meera Female Indian Detroit
Shaw Male Black Los Angeles
Shaw Male Black Detroit
Gloriana Female Hispanic Los Angeles
Gloriana Female Hispanic Detroit
Table 1: Examples of samples from the second subset of SilverSpoon.
By systematically pairing names with cities of varying economic statuses, we aim to assess how language models associate demographic attributes—such as race, gender, and location—with professions and other socioeconomic outcomes. This dataset provides a foundation for evaluating potential biases in the model’s predictions, allowing for a detailed examination of whether certain names and locations are more likely to be linked with lower- or higher-status professions, lifestyles, or narratives. We posit that this subset of our dataset can further serve as a synthetic dataset for many other applications in AI fairness research. For example, we believe this dataset will prove useful in testing fairness in social service tools, stereotype detection in AI systems, discrimination studies, and so on.
2.1.3 Part 3: Contextual Narrative Bias Analysis
To design this part of the dataset, we leveraged samples from the Demographic-Driven Profession Prediction dataset to generate short stories (Limit: 500 words) that explore the narratives surrounding individuals identified by their names and locations. Thus, each story was crafted using the unique combinations of the most common names associated with Black, White, Hispanic, and Indian origins, paired with the contexts of both, affluent and impoverished cities in the United States. This approach allows us to investigate how demographic attributes influence narrative construction, revealing underlying biases and assumptions embedded within the generated content.
By analyzing the stories produced, we aim to uncover patterns in how language models depict social status, challenges, and aspirations based on race, gender, and economic background. This narrative generation process serves as a powerful tool for examining implicit stereotypes, as the stories reflect not only the model’s understanding of individual identities but also the broader societal narratives associated with those identities.
@@ 3 Study Design
We experimented with a variety of open source and API based state-of-the-art large language models, namely, GPT-4o mini Bubeck et al. (2023), Llama3-8B Grattafiori et al. (2024), Gemma-7BTeam et al. (2024), Alpaca-13B Taori et al. (2023), Zephyr Tunstall et al. (2023), and Yi-6B 01.ai (2024). We choose to perform all prompting in a zero-shot manner because we are interested in examining the responses of models without any examples being provided to them. We hypothesize that even one example could introduce extraneous bias. Our experiments are run using A100s available on Google Colab and take a total of approximately 150 hours.
3.1 Normative Judgement Evaluation
Our main prompt asks the model to decide who is wrong in the given scenario, either "P1" or the "other person", for each of the 3000 samples in our dataset. Further, to perform high-caliber qualitative analysis, we elicit concise natural language explanations that justify the models’ answers. In particular, we find that the prompt used for GPT4 does not allow other models to engage with the text properly. Thus, we have a different prompt for the other models. We illustrate both prompts in Appendix K and describe the results of our prompts in Section 4.
3.2 Demographic Driven Narrative Bias Analysis
The prompt we design for this task asks the model to look at a sentence which reveals the name and location of a person, and make an educated guess as to that person’s profession. All models are instructed to answer in a few words at max, since this part of the methodology is designed to elicit any direct biases or stereotypes about certain demographic groups based on socioeconomic status.
3.3 Contextual Narrative Bias Analysis
We ask the model to leverage the information given about the a person (name and location only) and generate a story that is not more than 500 words. This word limit is set to facilitate compute while maintaining enough room for models to be creative and expressive. No other information or guidelines are provided here since the goal is to elicit indirect, subtle stereotypes or biases about certain groups based on social privilege.
@@ 4 Results
4.1 Normative Judgment Evaluation
We report per class accuracy, Cohen’s Kappa, Jaccard Similarity, Dice Similarity, and Matthew’s Correlation Coefficients against the upper-end labels and lower-end labels respectively. We do not use F1, precision, and recall in our evaluation since these metrics are calculated against a true-positive class, and our goal is to understand the similarity between model responses and each category of labels. We present multiple metrics in order to illustrate a complete picture of our quantitative analysis. Table 2 presents these results, To convert the model responses to a binary scale, we employ automatic evaluation, asking GPT4 to assess various responses and assign them to Class 1: Supporting the socioeconomically underprivileged, or Class 0: Otherwise. The prompt for this can be found in the appendix K.6.
@@ 4.1 Normative Judgment Evaluation
The range of values for accuracy, JSC, and DSC is 0 to 1 whereas the range for CKS and MCC is -1 to 1. For the latter two metrics, 0 indicates random chance agreement. We find that Alpaca has a negative CKC and MCC implying that its correlation with the lower-end labels is less than random, hence it agrees more with the upper class labels. AppendixJ offers a detailed discourse on the metrics of our choice and what they signify.
@@
We also perform an extensive qualitative analysis which is shown in AppendixD. This qualitative analysis reveals some interesting responses given by various models, and provides insight into the lack of empathy towards socioeconomically underprivileged groups.
4.2 Demographic Driven Narrative Bias Analysis
We first establish a baseline of the top 25 professions that make the most and least money, as officially reported by the U.S Department of Labor. For example, some professions which make the most amount of money are healthcare professionals (anesthesiologists, physicians, surgeons, oncologists, etc.), software engineers, computer hardware engineers, finance managers, marketing managers, and lawyers. Some professions which make the least amount of money are cooks, cashiers, fast food workers, waiters, maids, ushers, dishwashers and childcare providers. Then, we leverage Claude-SonnetAnthropic (2024) to assign a binary score of 0 or 1 to each model generated profession. If the profession is directly on the list of the richest or poorest professions, then assigning the score is simple. Otherwise, the model uses its training knowledge to assign the score and provides a one-line justification of the assigned score. The rationale of the model was manually empirically verified to be more accurate than 99%. The final step in our quantitative analysis pipeline is to group the names by gender and race to calculate percentages of rich and poor professions predicted by the model across locations.
We find that the models we test exhibit bias across multiple demographics. For example, white men are most commonly assigned professions such as "lawyer, doctor, software engineer", while white women are most commonly associated with professions such as "nurse, teacher, waitress". Furthermore, we find that the distribution of professions changes for men on an average of 7.4% across rich and poor locations and models, but by an average of 37.3% for women. This is demonstrated and quantified in Table3. The values in this table are an average of the values predicted by all models tested.
Gender Race City H.I Profession % L.I Profession %
Female White H.I 77.3% 22.7%
Male White H.I 88.5% 10.5%
Female White L.I 55.3% 44.7%
Male White L.I 60.3% 39.7%
Female Indian H.I 64.9% 35.1%
Male Indian H.I 90.2% 9.8%
Female Indian L.I 48.2% 51.8%
Male Indian L.I 57.6% 42.4%
Female Black H.I 39.1% 60.9%
Male Black H.I 38.7% 61.3%
Female Black L.I 38.5% 61.5%
Male Black L.I 41.3% 58.7%
Female Hispanic H.I 44.5% 55.5%
Male Hispanic H.I 47.7% 52.3%
Female Hispanic L.I 42.5% 57.5%
Male Hispanic L.I 46.4% 53.6%
Table 3: Model Demographic Driven Narrative Bias Assessment: Comparing Predictions Across Racial and Gender Groups in High and Low Income Locations. Values reported are an average of percentages predicted by all models tested. H.I and L.I stand for high income and low income, respectively.
Table3 is very revealing. We observe that in high-income cities, all models tested predict high-income professions much more frequently for white people and Indians,when compared to black people and Hispanic people. Further, in low-income cities, we see that people of color are more frequently associated with low-income professions. This is not true for white people. Another important observation is that across race and location, models always associate more higher paying jobs with males than females. This analysis quantitatively proves socioeconomic bias is interleaved with racial discrimination (especially with respect to the Hispanic and African American community), while also quantifying gender bias as an active problem in this area. An elaborate qualitative analysis reveals that socioeconomic bias against people of color is strongest in GPT4 and minimal in Yi-6b. This is consistent with the results obtained in the normative judgment evaluation.
One argument that could be made after viewing these results is that these models are a function of data and that their predictions may simply be a function of what the data collected from the real world may represent. However, we posit that this reasoning is not conducive to fair model development. While models may reflect real-life biases to some extent, their deployment can amplify and perpetuate those inequities in harmful ways. Models are not neutral—they are designed and trained with specific goals in mind, often without sufficient consideration of fairness. By excusing bias as a mere reflection of reality, we risk reinforcing systemic inequalities instead of challenging them. We firmly believe that responsible AI development involves identifying and mitigating biases, not just replicating them.
4.3 Contextual Narrative Bias Analysis
To analyze more subtle biases across socioeconomic classes, race, gender, and location, we utilize sentiment analysis and topic modeling.
We use sentiment analysis to analyze the language used in model responses (in this case, the generated stories) to prompts about different names and locations and assign a sentiment score to each response. The goal here is to see whether there is a pattern that can be observed across various demographic groups and locations. It is important to note that the prompts in this approach do not directly mention race or gender. We perform this analysis by utilizing a sentiment analysis model from HuggingFace which is a fine-tuned version of DistilBERT [^2] https://huggingface.co/tabularisai/robust-sentiment-analysis. This model looks at each generated story and assigns a label of either "very negative", "negative", "neutral", "positive", or "very positive". We further map these labels to a scoring system where -1 is assigned to labels "very negative" and "negative", 0 is assigned to label "neutral", and 1 is assigned to labels of "positive", "very positive".
Table4 presents a breakdown of the scores obtained. The values presented are an average across models tested, and the detailed model responses can be found in AppendixH.
Gender Race City Negative % Neutral % Positive %
Female White H.I 7.3% 50.2% 42.5%
Male White H.I 5.4% 42.7% 51.9%
Female White L.I 8.9% 60.1% 31.0%
Male White L.I 8.3% 59.6 32.1%
Female Indian H.I 7.2% 58.6% 34.2%
Male Indian H.I 6.6% 60.0% 33.3%
Female Indian L.I 9.1% 61.2% 29.7%
Male Indian L.I 8.9% 63.0% 24.1%
Female Black H.I 9.2% 66.7% 24.1%
Male Black H.I 9.3% 68.5% 22.2%
Female Black L.I 12.4% 69.0% 18.6%
Male Black L.I 11.7% 64.2% 24.2%
Female Hispanic H.I 10.3% 66.1% 23.6%
Male Hispanic H.I 10.8% 68.7% 20.5%
Female Hispanic L.I 11.8% 72.0% 16.2%
Male Hispanic L.I 12.3% 69.8% 17.9%
Table 4: Sentiment Analysis Across Demographics and Locations to Quantify Contextual Narrative Bias across all tested models. H.I and L.I stand for high income, and low income respectively.
The quantification of subtle biases also reveals a significant amount of insight. Perhaps one of the most notable observations is that groups that are traditionally considered minorities, such as Hispanic and African American (especially women) have a lower percentage of positive sentiment scores, and a higher percentage of negative sentiment scores. Furthermore, we observe that in locations that have high incomes, the percentage of positive sentiment score is the highest for privileged groups, and lowest for minority groups. Also, once again, we can see that for most races, men have a higher percentage of positive sentiment scores than women. This solidifies the building hypothesis that not only do large language models exhibit socioeconomic bias, but that this bias is further complicated by interleaved racial and gender bias.
We perform qualitative analysis of these results by performing topic modeling and getting the top 20 most frequent words for each group as represented in Table4. The results are demonstrated in AppendixI. To effectively summarize these results, consider that for the prompts that involve writing stories about white people in rich cities, some of the most common words used are "hard working, politician, lawyer, happy, intelligent" and some of the most common words used for Hispanic women are "struggling, beautiful, talented, hustler, and smart".